OpenStudy (anonymous):

serious simplifying complex fractions here!! 3 2 --- + --- x x+2v ------------------ 3 2 ---- - ---- x+2 x

OpenStudy (anonymous):

scratch that random v

OpenStudy (solomonzelman):

yes, I knew that !!

OpenStudy (solomonzelman):

\(\Huge\color{black}{ \frac{\frac{3}{x} + \frac{2}{x+2} }{\frac{3}{x+2} + \frac{2}{x} } }\) like this ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\Huge\color{black}{ \frac{\frac{3\color{darkgoldenrod}{\times (x+2)}}{x\color{darkgoldenrod}{\times (x+2)}} + \frac{2\color{darkgoldenrod}{\times x}}{(x+2)\color{darkgoldenrod}{\times x}} }{\frac{3\color{darkgoldenrod}{\times x}}{(x+2)\color{darkgoldenrod}{\times x}} + \frac{2\color{darkgoldenrod}{\times (x+2)}}{x\color{darkgoldenrod}{\times (x+2)}} } }\)

OpenStudy (solomonzelman):

can you tell me WHY am I doing this ?

OpenStudy (anonymous):

common denominator

OpenStudy (solomonzelman):

yes common denominator, tell me what you get after multiplying ad adding the fractions on top and bottom, please.

OpenStudy (anonymous):

\[5x+6 \over x(x+2)\]

OpenStudy (anonymous):

\[5x+4 \over x(x+2)\]

OpenStudy (solomonzelman):

Yes, the first one is the numerator, and the second is the denominator.

OpenStudy (solomonzelman):

\(\Huge\color{black}{ \frac{\frac{5x+6}{(x(x+2)} }{\frac{5x+4}{x(x+2)}} }\) is the same thing as, \(\LARGE\color{black}{\frac{5x+6}{(x(x+2)} \div\frac{5x+4}{x(x+2)} }\) and knowing that \(\LARGE\color{black}{\frac{A}{B} \div\frac{C}{D}=\frac{A}{B} \times\frac{D}{C} }\) can you tell me, what cancels out ?

OpenStudy (anonymous):

x(x+2)

OpenStudy (solomonzelman):

GOod ! And your answer is ?

OpenStudy (anonymous):

\[5x+6 \over 5x+4\]

OpenStudy (solomonzelman):

Yup :)

OpenStudy (anonymous):

thank you!

OpenStudy (solomonzelman):

Anytime !